x+y+z=4
2x-y+z=0
x-y+2z=-3

Answer by scott8148(6628)

adding 1st eqn and 2nd eqn ___ 3x + 2z = 4 ___ multiplying by 2 ___ 6x + 4z = 8

adding 1st eqn and 3rd eqn ___ 2x + 3z = 1 ___ multiplying by -3 ___ -6x - 9z = -3

adding new equations ___ -5z = 5 ___ z = -1

substituting ___ 2x + 3(-1) = 1 ___ x = 2

substituting ___ (2) + y + (-1) = 4 ___ y = 3